OpenStudy (anonymous):

Use prime factorization to find the LCM. a. 6, 16

OpenStudy (mertsj):

Can you write the prime factorization?

Parth (parthkohli):

I'd give you an example.

Parth (parthkohli):

OK, so if you have, say, \(7\) and \(21\), then first you do the prime factoring:\[7 = 7^1 \\ 21 = 7^1 \cdot 3^1\]The highest power of \(7\) in both numbers is \(1\) and for \(3\), it's \(1\). Simply multiply \(3^1\cdot 7^1 = \boxed{21}\).

OpenStudy (anonymous):

that's hoe it came. so i don't know how to do it. i am not good at math at all

Parth (parthkohli):

Can you do the prime factorization?

OpenStudy (anonymous):

i don't even know that that is

Parth (parthkohli):

Every number can be expressed as a product of prime numbers. So there's a way that you can express \(6\) as the product of some prime numbers. The same goes for \(16\).

OpenStudy (anonymous):

ok

Parth (parthkohli):

So try to express \(6\) and \(16\) as product of primes?

OpenStudy (anonymous):

would it be 48?

Parth (parthkohli):

How did you get \(48\)?

OpenStudy (anonymous):

i don't know i just picked a number

Parth (parthkohli):

That is incorrect.

OpenStudy (skullpatrol):

Prime factors of positive integers can be found by using the primes in order as divisors.

OpenStudy (skullpatrol):

The prime factorization of a positive integer is the expression of the integer as a product of prime factors.

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):

np